A car moving with a speed of 40 km/h can be stopped by applying brakes at least after 2m. If the same car is moving with a speed of 80 km/h, what is the minimum stopping distance?

(a)8m

(b)6m

(c)4m

(d)2m

Grade:10

4 Answers

Jitender Pal
10 years ago

(a)
If F is retarding force and s the stopping distance then
    1/2  m2   =Fs
    For same retarding force, s a v2
        s_2/s_1 = (v_2/v_1 )^2  =((80 km/h)/(40 km/h))^2  =4    
       s2 = 4s1 = 4 × 2 = 8m

vivek sharma
6 years ago
F is retarding force, and x is stopping distance. Then F. x = 1/2 mv² F. 2 =1/2m(40)² F=1/4.m.1600 X’ = mv²/2F X’ = 6400m/1600m = 8 metre
Nazir Hossain
6 years ago
First calculate the acceration in first case v ^2= 2aSa=v^2/2s , a= 1600/2*2, a=400m/s^2Now, calculate distance in 2nd caseS= v^2/2a , S= 6400/2*400, S= 8 m
ammalu
6 years ago
we known that v2=u2+2as then a=u2/2s =(50* 5/ 18)2 /2*6
s=u2/2a then (100*5/18)       =24
                 2 * (50*5/ 18)2
                          12